OpenStudy (anonymous):

how do you solve 2 cos(x + 30) cos(x - 30) = 1 with restriction x E [-180, 180]?

OpenStudy (ranga):

Use the identities cos(A+B) = cos(A)cos(B) - sin(A)sin(B) and cos(A-B) = cos(A)cos(B) + sin(A)sin(B) sin(30) = 1/2 cos(30) = sqrt(3)/2 Substitute and simplify

OpenStudy (anonymous):

thank you!

OpenStudy (ranga):

you are welcome.

OpenStudy (ranga):

What are you getting as a solution for x?